OpenStudy (anonymous):

Solve –4(x + 3) = 16

OpenStudy (anonymous):

Distribute the -4. \[a(b+c) = (a\times b) + (a \times c)\]

jhonyy9 (jhonyy9):

-4x-12=16 -4x=28 x=28/-4 x=-7